luz paid P1,080 in interest on a P3,060 loan she got from the credit union. The term of the

loan is 3 years. Find the rate.


1
Expert's answer
2021-12-12T18:17:44-0500

1) 3 years = 36 months.


2) P1,080 + P3,060 = P4,140 (total payment).


3) "{\dfrac{1,080}{4,140}} \approx 0.26"


4) Since P0.26 is 26%, then 26% is a total rate.


5) "{\dfrac{0.26}{36}} \approx 0.72"% - is a month rate.


Answer: 0.72%